# Erasing ‘Ugly’ from the Internet: Propagation of the Beauty Myth in Text-Image Models

### v1

@@ 5.2 Hypothesis 2
Figure  4 shows that participants selected SFW for $55.9\%$ of the generated images, and $12.1\%$ for the closest next category and then $10\%$, indicating that overall, $78\%$ of the images were not rated as sexualised or did not contain nudity. However, the presence of any sexualised imagery is notable. We chose the F-Lite model given the emphasis that it has been trained exclusively on images marked SFW. However, as we’ve discussed in the previous section, how sexualised an image appears depends on one’s lived experience, and we lack transparency regarding the quality of the SFW labels given to the 80 million images Ryu et al. (2025) used to train the model. Figure 5 shows some examples of nudity (censored) generated by both models.
@@
Stark results for $Q3$ are depicted in Figure  4, which shows that whiteness is overwhelmingly the skin tone generated in the images, with $86.5\%$ of the images rated by participants as within Fitzpatrick scales of I-III ($45.2\%$ images rated as skin type I, $28.2\%$ as II, and $13.1\%$ as III), consistent with Buolamwini & Gebru (2018). For the two-way ANOVA analysis, both gender ($F=3.544,p_{adj}=.044$) and polarity ($F=12.706,p_{adj}<.001$) had significant main effects. However, the interaction between the two conditions was not significant ($F=2.198,p_{adj}=.089$). An exploratory Tukey post-hoc comparison ($\alpha=0.5$) between group combinations suggest that generated images of women using neutral prompts consistently have the lightest skin tones, significantly more than images generated of non-binary people stemming from positive and negative prompts ($p_{adj}<.001$ and $p_{adj}=.002$ respectively) and images of men generated using positive prompts ($p_{adj}=.001$). We interpret these results cautiously, given that this may not reflect a true interaction between the groups.
@@ 6 Downstream harms and consequences
As shown in 2, we already see the damaging effects that social media has on people, provoking body dysmorphic disorder and general dissatisfaction with one’s own appearance Laughter et al. (2023); Ateq et al. (2024); Thawanyarat et al. (2022); Laughter et al. (2023); Maymone & Kroumpouzos (2022). In this section, we discuss the answers to our research question in terms of the possible downstream harms that beauty ideals encoded in generative AI could have on society.
@@ 6.1 Erasure of features and pollution of the data stream
While results for $H_{1}$ were only partially supported, we still qualitatively observe this phenomenon occurring in our generated images (for e.g. in Figure 5 (b), the models’ apparent inability to generate a woman with narrow lips). Additionally, we see clear results in $H_{2}$ in terms of younger-looking individuals, and Fitzpatrick skin tones that are fairer. We additionally discussed our results that may indicate a bias between perceived youth with sexualisation. We know for example that facial recognition datasets are heavily biased towards lighter-skinned individuals Buolamwini & Gebru (2018), that researchers have found a correlation between sexually explicit content and hateful language Birhane et al. (2023), and that an ‘attractive ideal’ in generative image models produces faces that approximate a ‘white ideal’ Bianchi et al. (2023).
However, what happens when the volume of these images generated far exceeds the amount of content created by humans? At scale, such patterns may shape public perception, particularly for vulnerable populations, e.g. teenagers with skin conditions or older people that naturally develop wrinkles. As Wolf (1991) shows, society already enforces rigid beauty standards upon us, and these are now being amplified by generative AI Bianchi et al. (2023), rather than using this technology to challenge these norms. Furthermore, this phenomenon could lead to the pollution of data streams. For example, images of female breasts created by generative AI that are often oversized and sexualised Kenig et al. (2023), could swamp carefully curated medical images of healthy breasts or those used for cancer screening.
@@ 6.2 Negative prompting
It is essential to additionally address the process of creating the actual images, which often involves negative prompting (not to be confused with prompts that we use encoded with polarity). This method, popularised in recent image generation models, in particular Stable Diffusion Ban et al. (2024), involves providing ‘undesirable’ qualities for image generation, i.e. what not to generate. While we did not use negative prompting in our pipelines (it is not recommended for this particular version of Stable Diffusion fofr (2024)), Freepik suggests it for the F-Lite model [^13] https://github.com/fal-ai/f-lite/blob/main/NEGATIVE_PROMPT.md as of 21st May 2025.. While some words to include in the negative prompt serve the purpose of not generating human bodies with four legs, some other words are concerning; such as “[…] distorted proportions, deformed, bad anatomy, extra limbs, missing limbs, disfigured face, poorly drawn hands, poorly rendered face, mutation […]”. In other extreme cases, “[…] ugly, …morbid, mutilated, …, dehydrated, bad anatomy, bad proportions, …, disfigured, gross proportions, malformed limbs […]”. [^14] As given for this ‘realistic’ image generation model available here https://huggingface.co/SG161222/Realistic_Vision_V5.1_noVAE as of 21st May 2025. Further analysis is warranted to see how much our prompts encoded with negative polarity overlap with words provided in negative prompting guides. According to Desai & Vasconcelos (2024), the word ‘ugly’ is a consistently used negative prompt. Synthetically created images are (over)saturating the internet landscape. Developers of models are encouraging people to codify ‘desirable’ qualities in these images (via negative prompting), and remove what is seen as ‘undesirable’. Via negative prompting alone, the title “Erasing ‘Ugly’ from the Internet’ seems fitting; however who gets to decide what is ‘ugly’, which proportions are bad, what is ‘undesirable’? Negative prompting shows that beauty ideals are not only present as biases in image models, but are actively perpetuated by the model developers.

### v2

@@ 5.2 Hypothesis 2
Figure  4 shows that participants selected SFW for $55.9\%$ of the generated images, and $12.1\%$ for the closest next category and then $10\%$, indicating that overall, $78\%$ of the images were not rated as sexualised or did not contain nudity. However, the presence of any sexualised imagery is notable. We chose the F-Lite model given the emphasis that it has been trained exclusively on images marked SFW. However, as we’ve discussed in the previous section, how sexualised an image appears depends on one’s lived experience, and we lack transparency regarding the quality of the SFW labels given to the 80 million images (ryu2025flite) used to train the model. Figure 5 shows some examples of nudity (censored) generated by both models.
@@
Stark results for $Q3$ are depicted in Figure  4, which shows that whiteness is overwhelmingly the skin tone generated in the images, with $86.5\%$ of the images rated by participants as within Fitzpatrick scales of I-III ($45.2\%$ images rated as skin type I, $28.2\%$ as II, and $13.1\%$ as III), consistent with (buolamwini-gebru-2018-gender). For the two-way ANOVA analysis, both gender ($F=3.544,p_{adj}=.044$) and polarity ($F=12.706,p_{adj}<.001$) had significant main effects. However, the interaction between the two conditions was not significant ($F=2.198,p_{adj}=.089$). An exploratory Tukey post-hoc comparison ($\alpha=0.5$) between group combinations suggest that generated images of women using neutral prompts consistently have the lightest skin tones, significantly more than images generated of non-binary people stemming from positive and negative prompts ($p_{adj}<.001$ and $p_{adj}=.002$ respectively) and images of men generated using positive prompts ($p_{adj}=.001$). We interpret these results cautiously, given that this may not reflect a true interaction between the groups.
@@ 6 Downstream harms and consequences
As shown in 2, we already see the damaging effects that social media has on people, provoking body dysmorphic disorder and general dissatisfaction with one’s own appearance (LAUGHTER202328; ateq_2024; thawanyarat2022zoom; LAUGHTER202328; MAYMONE2022554). In this section, we discuss the answers to our research question in terms of the possible downstream harms that beauty ideals encoded in generative AI could have on society.
@@ 6.1 Erasure of features and pollution of the data stream
While results for $H_{1}$ were only partially supported, we still qualitatively observe this phenomenon occurring in our generated images (for e.g. in Figure 5 (b), the models’ apparent inability to generate a woman with narrow lips). Additionally, we see clear results in $H_{2}$ in terms of younger-looking individuals, and Fitzpatrick skin tones that are fairer. We additionally discussed our results that may indicate a bias between perceived youth with sexualisation. We know for example that facial recognition datasets are heavily biased towards lighter-skinned individuals (buolamwini-gebru-2018-gender), that researchers have found a correlation between sexually explicit content and hateful language (birhane2023laions), and that an ‘attractive ideal’ in generative image models produces faces that approximate a ‘white ideal’ (bianchi-etal-2023-easily).
However, what happens when the volume of these images generated far exceeds the amount of content created by humans? At scale, such patterns may shape public perception, particularly for vulnerable populations, e.g. teenagers with skin conditions or older people that naturally develop wrinkles. As wolf1991beauty shows, society already enforces rigid beauty standards upon us, and these are now being amplified by generative AI (bianchi-etal-2023-easily), rather than using this technology to challenge these norms. Furthermore, this phenomenon could lead to the pollution of data streams. For example, images of female breasts created by generative AI that are often oversized and sexualised (kenig-etal-2023-human), could swamp carefully curated medical images of healthy breasts or those used for cancer screening.
@@ 6.2 Negative prompting
It is essential to additionally address the process of creating the actual images, which often involves negative prompting (not to be confused with prompts that we use encoded with polarity). This method, popularised in recent image generation models, in particular Stable Diffusion (ban2024understanding), involves providing ‘undesirable’ qualities for image generation, i.e. what not to generate. While we did not use negative prompting in our pipelines (it is not recommended for this particular version of Stable Diffusion (fofr_2024)), Freepik suggests it for the F-Lite model [^13] https://github.com/fal-ai/f-lite/blob/main/NEGATIVE_PROMPT.md as of 21st May 2025.. While some words to include in the negative prompt serve the purpose of not generating human bodies with four legs, some other words are concerning; such as “[…] distorted proportions, deformed, bad anatomy, extra limbs, missing limbs, disfigured face, poorly drawn hands, poorly rendered face, mutation […]”. In other extreme cases, “[…] ugly, …morbid, mutilated, …, dehydrated, bad anatomy, bad proportions, …, disfigured, gross proportions, malformed limbs […]”. [^14] As given for this ‘realistic’ image generation model available here https://huggingface.co/SG161222/Realistic_Vision_V5.1_noVAE as of 21st May 2025. Further analysis is warranted to see how much our prompts encoded with negative polarity overlap with words provided in negative prompting guides. According to desai2024improving, the word ‘ugly’ is a consistently used negative prompt. Synthetically created images are (over)saturating the internet landscape. Developers of models are encouraging people to codify ‘desirable’ qualities in these images (via negative prompting), and remove what is seen as ‘undesirable’. Via negative prompting alone, the title “Erasing ‘Ugly’ from the Internet’ seems fitting; however who gets to decide what is ‘ugly’, which proportions are bad, what is ‘undesirable’? Negative prompting shows that beauty ideals are not only present as biases in image models, but are actively perpetuated by the model developers.